OpenStudy (anonymous):

Solve the equation. 4x+3+6x+9=9x+5

OpenStudy (anonymous):

Collect the Like Terms

OpenStudy (anonymous):

can you show me an example?

OpenStudy (anonymous):

4x+6x-9x=5-3-9

OpenStudy (anonymous):

x=-7

OpenStudy (anonymous):

like, you would do 4x+6x+9x then add the rest of the numbers...get it?

OpenStudy (anonymous):

so would it be 19x? @ChemicalX

OpenStudy (anonymous):

Yup. :) Now you do the same with the others.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

no 10x-9x=x

OpenStudy (anonymous):

not 19x

OpenStudy (anonymous):

that's not what i got...